The point(-6,-6) lies on a circle whose center is at (-5,4) . Find an equation of this circle.

Answer by scott8148(6628)

for a circle centered at (h,k) with radius r ___ (x -h)^2 + (y - k)^2 = r^2

in this case ___ (x + 5)^2 + (y - 4)^2 = 101